OpenStudy (anonymous):

What is the slope of the line 5x + 2y = 10 Answer none of the above

OpenStudy (ash2326):

We have \[5x+2y=10\] Standard line equation in slope intercept form \[y=mx+c\] \[\text{m is the slope}\] \[\text{ c is the y intercept}\] Let's convert our equation in this form \[5x+2y=10\] \[2y=10-5x\] \[y=-\frac{5}{2}x+5\] Now tell me what's the slope?